unicornquenn71:

Rewrite 2 2/5 as an improper fraction. 5/9 12/5 9/2 9/5

unicornquenn71:

i think it is 12/5

makaylab:

Simplify 2 and 2 over 52 2 5 = 12 over 5 12 5 Step 1 of 1: Simplify, sub-step a: Reduce fraction to lowest terms. Reduce fraction to lowest terms 1 is the greatest common divisor of 2 and 5. The result can't be further reduced.

Angelwolf:

Step 1 Multiply the denominator by the whole number 5 × 2 = 10 Step 2 Add the answer from Step 1 to the numerator 10 + 2 = 12 Step 3 Write answer from Step 2 over the denominator 12/5 = 12 5

makaylab:

yep thats righ angel wolf and unicornquenn71

unicornquenn71:

ok

unicornquenn71:

on to the next one